OpenStudy (anonymous):

√-75 write the complex number in stardard form

OpenStudy (shadowfiend):

So, the same applies as with your previous question. Try coming up with the answer based on that and post it, then we'll see if it's right, and if it's wrong I'll explain the process again.

OpenStudy (anonymous):

i know can you example the reason again

OpenStudy (shadowfiend):

So if you have any negative number -a, that's the same as saying -1 * a. Above, we have -75. That's the same as -1 * 75. Finally, \(\sqrt{-1} = i\).

OpenStudy (anonymous):

so thats the answer?^

OpenStudy (shadowfiend):

Not quite. Since we know -1 * 75 = -75, we get: \[\sqrt{-1 * 75}\] \[\sqrt{-1}\sqrt{75}\] What's \(\sqrt{-1}\) again?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

i

OpenStudy (anonymous):

isn't it\[5\sqrt{3}i\]

OpenStudy (shadowfiend):

Yep, that's part 2, factoring out the 75 into 25 * 3 and doing that goodness :)